OpenStudy (catpidigm):

1. Find the 15th term in the Arithmetic Sequence An = 3n-7 2. Find the 10th term in the Arithmetic Sequence -2, -5, -8, -11 Please help!

OpenStudy (mhchen):

1. You are given the explicit formula of An = 3n - 7 So to find the 15th term, you plug it in: A(15) = 3(15)-7 A(15) = 45 - 7 A(15) = 38 <-- this is how you determine the term. 2. You are given an arithmetic sequence, so you should make a formula based off of it. If you find the difference between each term, it's -3 because -2 - 3 = -5 -5 - 3 = -8 -8 - 3 = -11 So the -3 would be the SLOPE Now how do we find the constant? We know that -2 is the first term so -2 = -3(1) + x <-- plugged 1 in because (-2) is the FIRST TERM where x is the constant we're looking for so -2 = -3 + x 1 = x So our constant is 1. Now we have our arithmetic explicit formula: An = -3(n) + 1 Using what we've done for problem one, can you figure out the 10th term using this formula we've made?

OpenStudy (catpidigm):

Ok one second, kinda slow at math... btw thank you so much! fan and medal :)

OpenStudy (catpidigm):

is -29 the answer to question 22?

OpenStudy (catpidigm):

There are 4 questions I need help with. Can you help me with them please? (Medal)

OpenStudy (mhchen):

Yeah, I got -29 as well. Sorry I g2g though. I can't help with anymore problems. Too tired :\

OpenStudy (catpidigm):

Ok :) But thank you so much!

OpenStudy (catpidigm):

Going to close this question and ask another with questions 3 and 4.